Week of 160307

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (Chair and minutes), Doug Benjamin (ATLAS), Asa Hsu (ASGC), Kate and Liana Lupsa (DB), Belinda and Jesus Lopez (Storage), Ben Jones (Computing), Maria Dimou (GGUS), Eric Vaandering (CMS)
  • remote: Michael Ernst (BNL), Mateo (CNAF), David Mason (FNAL), Rolf Rumler (IN2P3), Sang Un Ahn (KISTI), Dimitri (KIT), Onno Zweers (NL-T1), Kyle (OSG), Pepe Flix (PIC), John Kelly (RAL), Di Qing (TRIUMF)

Experiments round table:

  • ATLAS reports (raw view) -
    • Amazon EC2 scale test postponed from last week to be continued this week
    • Taiwan site requests until they solve issues with their CVMFS stratum one that it is not used
    • shortage of tasks caused some sites to be empty of ATLAS jobs at the end of the week

Maria explains that ATLAS and ASGC have been already in touch to decommission CVMFS Stratum 1 for the time being.

  • CMS reports (raw view) -
    • Nagios issue that broke HTCondor submission to CREAM CEs
      • Recovered just fine after things were fixed
    • Report of vocms0114 (CRAB3 machine) with a high rate of DNS requests
      • No complaint since initial Friday night complaint - possibly related to FNAL downtime?

Eric asks whether there are more news about the CRAB3 machine. Maria asks whether there is any ticket opened and Eric says there was only one mail sent. This will be followed up offline.

  • ALICE -
    • Team ticket GGUS:119885 for CERN Wed morning because of 20k+ ghost jobs in LSF.
      • Fixed Wed afternoon.
    • New record reached Thu morning: 97310 running jobs!
      • Also thanks to 8.5k running on the HLT farm!
    • Ticket GGUS:119941 for SAM Thu evening because of empty dashboard pages after DB intervention.
      • Fixed Fri morning.
    • Fri morning EOS at CERN needed to be restarted to cure a problem (OTG:0029122).
      • This caused many job failures around the grid for ~2h.
    • A flood of requests for users to re-sign the AUP led to a support load on VO admins.
      • A randomization proposal is under discussion.

  • LHCb reports (raw view) -
    • Site Issues
      • T0: HTCondor-CE for job submission working. Pilot log retrieval not operational yet.
      • T1:
        • IN2P3: Pilots failing b/c exceeding memory limits (GGUS:119886). Problematic event type for the currently ongoing validation of the next version of the simulation framework.

IN2P3 asks whether the ticket could be reopened so they can follow up. Stephan clarifies that the issue is on LHCb side so it's understood this is not IN2P3 problem.

Sites / Services round table:

  • ASGC: There was a NFS server downtime during the weekend. Recovery still ongoing.
  • BNL: NTR
  • CNAF: NTR
  • FNAL: FNAL sitewide power outage as of 22h00 FNAL time (about 5h00 AM CERN time) Recovered without problems.
  • GridPP: NA
  • IN2P3: Most of the sites services will be in downtime all day on Tuesday, March 15th. In particular, the mass storage system, iRods, dCache and the batch system will be unavailable. For the latter one we will change its version which makes necessary to disallow submission from the late evening of the 14th onward and also to suppress all remaining jobs in queue in the morning of the 15th.
  • JINR: NA
  • KISTI: NTR
  • KIT: Postponed tape downtime scheduled for this week.
  • NDGF: NA
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: NTR
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • CREAM and HTCondor CEs, LSF and ARGUS down for NFS filer migration (GOCDB) on Tuesday 15.03 from 10h00 to 12h00. No batch submission possible on LSF. Running jobs not affected.
    • MyProxy down for NFS file migration (GOCDB) on Thursday 17.03 from 10h00 to 12h00.
  • CERN storage services:
    • CASTORALICE was upgraded to the latest release (2.1.16-0). Other LHC instances will be upgraded in due course (before the LHC restart).
  • CERN databases:
    • An OS update in the Connection Manager for Atlas and CMS will take place tomorrow at 2pm. During the intervention it will not possible to connect to the following databases from the outside: ATLARC, intr, int8r, int2r, int9r
  • GGUS: NTR
  • MW Officer: NTR

AOB:

  • REMINDER: T1s input on 2016 pledges at the next Ops Coord on the 17th March (Indico Agenda).
Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2016-03-07 - KateDziedziniewicz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback